what is the ordered pair for 2x+4y=4 and 5x+3y=-16

Answer by stanbon(75887)

what is the ordered pair for 2x+4y=4 and 5x+3y=-16
---------
Simplify::
x + 2y = 2
5x + 3y = -16
-----
Modify for elimination:
5x + 10y = 10
5x + 3y = -16
-------------------
Subtract and solve for "y"::
7y = 26
y = 26/7
-------
Solve for "x"::
x = -2y + 2
x = -2(26/7) + (14/7)
---
x = -36/7
----------------
Cheers,
Stan H.
---------